Operations Report – March 27th

Crew 191 Operations Report 27/03/2018

SOL:3

Name of person filing report: Wataru Okamoto

Non-nominal systems: N/A

Notes on non-nominal systems: N/A

Generator (hours run):13hours

Solar SOC 96% (Before generator is run at night)

Diesel 25%

Propane 70psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) ? 7gallons.

Water (trailer) 60%

Water (static) 80%

Trailer to Static Pump used ? no

Water (loft) Static to Loft Pump used yes

Water Meter: 7:50 6 132778.1->pump 14 16:00 10 132790.2

Toilet tank emptied: yes

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) –

Oil Added –

ATV Fuel Used: – 350.2 45%->75%

Hours the ATVs were Used today: –

Notes on ATVs:-

Deimos rover used: –
Hours:
Beginning charge:
Ending charge:
Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: –
Hours:27.9h
Beginning charge:100%
Ending charge: 100%
Currently charging: 100%

Opportunity rover used: –
Hours: 26.9h
Beginning charge: 100%
Ending charge: 100%
Currently charging: 100%

Curiosity rover used: –
Hours:22.1h
Beginning charge: 100%
Ending charge: 100%
Currently charging: 100%

HabCar used and why, where? –

General notes and comments: -I didn’t turn off the Diesel generator in this morning. I turned off the Diesel generator after 50 minuites. It’s my mistake.I’ll turn off when amount of battery is over 80%.

Summary of internet: Under control

Summary of suits and radios:No.5 Backpack Battery was 6.8V after charging. cSo another 13V battery was installed.2 Backpack belt have repaired

Summary of Hab operations:No problem

Summary of GreenHab operations: No problem

Summary of ScienceDome operations: Only use checking the solar system.

Summary of RAMM operations: N/A

Summary of health and safety issues: No problem

Questions, concerns and requests to Mission Support:Please charge the Diesel gas.

Greenhab Repot – March 27th

GreenHab Report
Kai Takeda
27 MAR 2018

Environmental control:

Ambient with window/door open
Shade cloth on
Fan on
Heater on(morning)

Average temperatures: 22.8℃
Low 18.5℃
High 31.3℃

Hours of supplemental light: None

Changes to crops: Added hydroponic crops.

Daily water usage for crops: 2gal

Time(s) of watering for crops: 08:30 and 15:30

Moringa research observations: None

Changes to research plants: No

Aquaponics: None

Narrative: I set a Japanese hydroponic machine “foop” and my hydroponic crops.
I harvested tomatoes, green peppers, leaf lettuce and herb.

Support/supplies needed: None

EVA Report – March 27th

EVA Report

EVA #3 – 27 Mar 2018 – Sol3
Crew members: Fumiei Morisawa (EVA leader), Venzha Christ, Makoto Kawamura

Objectives:
– Rubbish collections around MDRS
– Lecture about GPS device handling
– Scouting for Venzha’s CMEs project

Operations:
9:00 – Entering the airlock
9:03 – Collecting an engineering log in EVA
9:34 – Start to garbage collecting and scouting
10:50 – Turn back to the HAB
11:26 – Enter the airlock
11:29 – EVA end

Narrative:
It is sunny and wind is calm today, a condition is nice for collecting rubbish. We have patrolled the foot of North Ridge and collected various rubbish. A piece of cloth, plastic beverage bottle, old glove, and so on. Some of the large object that we couldn’t carry by hand are leaved there as it was. So we need to recollect them on another EVA with a vehicles later on. All the rubbish have recorded by GPS coordinates and photos data.

Vehicles
– none